OpenStudy (anonymous):

if one of the roots of the equation 2x2+px+9=0 is twice the other root. find the possible value of p

OpenStudy (anonymous):

p=-9 and roots are 1.5 and 3

OpenStudy (anonymous):

can u show me the method?

OpenStudy (anonymous):

method- the way u do it~

OpenStudy (anonymous):

let on root be y, so the other one is 2y. so (x-2y)(x-y)=0 x^2-3xy+y^2=0 is we multiply the eqn by 2 2x^2-6xy+2y^2=0 Now equate the coefficients.. u will get the answers.

OpenStudy (anonymous):

owh..ok :D tq so much~

OpenStudy (anonymous):

Please rate the answer if you understand..